INSERT A RATIONAL NUMBER BETWEEN 2/9 AND 3/8, AND ARRANGE IN DESCENDING ORDER?

Grade:

2 Answers

Abhishek Kumar
9 years ago
Take LCM of 8 and 9 and make their denominators same.
2/9 = 16/72
3/8 = 27/72
All numbers,
17/72, 18/72, 19/72, 20/72......26/72 are in between the two numbers.
Saloni Rakholiya
9 years ago
Take the Lcm ….make the denominators same (equivalent fraction) and then u can easily arrange it